# Mode-conditioned music learning and composition: a spiking neural network inspired by neuroscience and psychology

### v1

@@ 3.2 Music Data Encoding
The external stimuli(such as keys, pitches,etc.) are transformed into the input current $I$ that is suitable for the computational neuron model to receive. For example, a G-Major musical piece activates the neurons in the major cluster in the first layer and G major in the second layer. Equation.3-.4 describes the transformation for the mode and key clusters, respectively.
@@
In this study, we use the Izhikevich neural model [43] to simulate the behaviors of neurons within both the musical theory subsystem and the sequential memory system. The model is described as follow equations:
@@ 3.3.1 Synaptic Creation
Neuroscientific research has demonstrated that neural electrical activities and the dynamics of axonal growth cone movements are interdependent and coordinated. Electrical activities serve as feedback signals for the navigation of growth cones, while the movement of growth cones and the formation of new synapses subsequently influence the electrical activities within the neural network. [44, 45, 46]. In this study, we simplify this complex process by establishing rules for the formation of new synaptic connections and neural circuits in response to the input of musical information.
@@
The establishment of novel neural pathways mainly occurs between the musical theory subsystem and the sequential memory subsystem, which indicates that the collaborative learning and interaction of these two subsystems. Figure. 7 illustrates a simple example of how the model learns a four-part music piece in A minor. The model starts by receiving a set of notes $N_{1}={\{N_{1,1},N_{1,2},N_{1,3},N_{1,4}\}}$ from four parts, accompanied by their respective symbolic pitches $X_{1}=\{{76,69,60,45\}}$. These notes fire the corresponding neurons in the mode and key clusters (highlighted by red parallelograms and circles) according to Eq.3-4 and Eq.7. Simultaneously, the notes $N_{1}$ also prompt the activation of neurons(represented by red circles) situated across various minicolumns within the pitch subnetworks of the sequential memory subsystem. These neural oscillations facilitate the formation of both feedforward and feedback connections(marked as red double arrows) between neurons in the musical theory subsystem and those in the sequential memory system by the rule 8. The synaptic weights are initially set at random values, reflecting the stochastic nature of early learning stages. These weights will be subsequently updated by STDP learning rule as the learning process unfolds. The transmission delay $del$ for these new connections is set to zero in this phase, implying instantaneous signaling for simplicity. It is worth noting that the prior knowledge relationships, embodied as synaptic connections, are already established and depicted by blue arrows. Upon receiving the subsequent note set $N_{2}$, a similar pattern ensues, triggering the activation of corresponding neurons across both subsystems. When no connections link two oscillating neurons, novel connections are established spontaneously. This intricate process is characterized by the creation of neural circuits spanning the two subsystems.
@@ 3.3.2 Synaptic Plasticity
This paper utilize the modified STDP(Spike-Timing Dependence Plasticity)[47] learning rule to account for the spiking transmission delay when updating synaptic weights during the learning process, as shown in Eq. 10.
@@
$I_{ij}^{P}(t)$ denotes the total input of the $i$-th neuron in $j$-th minicolumn located in the pitch subnetwork at time step $t$. $I^{P}_{E\_ij(t)}$ is the current caused by the external stimuli(pitch value), $w^{MTS}_{s}$ means the $s$-th synaptic weight from the music theory subsystem, and $w^{P}_{ri}$ represents the $r$-th synaptic weight from the neurons in pitch subnetwork. The detailed discussion regarding this compounent can be found in our earlier work[38]. Similarly, the input of neurons in the duration subnetwork can be expressed by Eq. 12. A notable difference, however, is that neurons in duration subnetwork do not receive the signals from music theory subsystems.
@@ Acknowledgments
Author Contributions
Qian.L designed the study, implemented the code, conducted the experiments, and drafted the manuscript. Y.Z. supervised the project and contributed to the manuscript writing. MHR.T. assisted with code maintenance.
Declaration of interests

### v2

@@ 3.2 Music Data Encoding
The external stimuli(such as keys, pitches,etc.) are transformed into the input current $I$ that is suitable for the computational neuron model to receive. For example, a G-Major musical piece activates the neurons in the major cluster in the first layer and G major in the second layer. Equation.1-.2 describes the transformation for the mode and key clusters, respectively.
@@
In this study, we use the Izhikevich neural model Izhikevich (2003) to simulate the behaviors of neurons within both the musical theory subsystem and the sequential memory system. The model is described as follow equations:
@@ 3.3.1 Synaptic Creation
Neuroscientific research has demonstrated that neural electrical activities and the dynamics of axonal growth cone movements are interdependent and coordinated. Electrical activities serve as feedback signals for the navigation of growth cones, while the movement of growth cones and the formation of new synapses subsequently influence the electrical activities within the neural network. Spitzer (2006); Hua and Smith (2004); Zheng and ming Poo (2007). In this study, we simplify this complex process by establishing rules for the formation of new synaptic connections and neural circuits in response to the input of musical information.
@@
The establishment of novel neural pathways mainly occurs between the musical theory subsystem and the sequential memory subsystem, which indicates that the collaborative learning and interaction of these two subsystems. Figure. 3 illustrates a simple example of how the model learns a four-part music piece in A minor. The model starts by receiving a set of notes $N_{1}={\{N_{1,1},N_{1,2},N_{1,3},N_{1,4}\}}$ from four parts, accompanied by their respective symbolic pitches $X_{1}=\{{76,69,60,45\}}$. These notes fire the corresponding neurons in the mode and key clusters (highlighted by red parallelograms and circles) according to Eq.1-2 and Eq.5. Simultaneously, the notes $N_{1}$ also prompt the activation of neurons(represented by red circles) situated across various minicolumns within the pitch subnetworks of the sequential memory subsystem. These neural oscillations facilitate the formation of both feedforward and feedback connections(marked as red double arrows) between neurons in the musical theory subsystem and those in the sequential memory system by the rule 6. The synaptic weights are initially set at random values, reflecting the stochastic nature of early learning stages. These weights will be subsequently updated by STDP learning rule as the learning process unfolds. The transmission delay $del$ for these new connections is set to zero in this phase, implying instantaneous signaling for simplicity. It is worth noting that the prior knowledge relationships, embodied as synaptic connections, are already established and depicted by blue arrows. Upon receiving the subsequent note set $N_{2}$, a similar pattern ensues, triggering the activation of corresponding neurons across both subsystems. When no connections link two oscillating neurons, novel connections are established spontaneously. This intricate process is characterized by the creation of neural circuits spanning the two subsystems.
@@ 3.3.2 Synaptic Plasticity
This paper utilize the modified STDP(Spike-Timing Dependence Plasticity)Bi and Poo (1998) learning rule to account for the spiking transmission delay when updating synaptic weights during the learning process, as shown in Eq. 8.
@@
$I_{ij}^{P}(t)$ denotes the total input of the $i$-th neuron in $j$-th minicolumn located in the pitch subnetwork at time step $t$. $I^{P}_{E\_ij(t)}$ is the current caused by the external stimuli(pitch value), $w^{MTS}_{s}$ means the $s$-th synaptic weight from the music theory subsystem, and $w^{P}_{ri}$ represents the $r$-th synaptic weight from the neurons in pitch subnetwork. The detailed discussion regarding this compounent can be found in our earlier work Liang et al. (2020). Similarly, the input of neurons in the duration subnetwork can be expressed by Eq. 10. A notable difference, however, is that neurons in duration subnetwork do not receive the signals from music theory subsystems.
@@ 3.4 Mode-Conditioned Music Generation
4 Datasets
4.1 Dataset Description
J.S. Bach’s four-part chorales(Bach): To make the model ’listen’ more music works, we employ the J.S. Bach’s four-part chorales dataset, which is available through the Music21 python package Cuthbert and Ariza (2010). It includes 408 famous chorales in musicxml format, with 219 in major keys and 189 in minor keys, excluding Db and Gb major, as well as Db, Eb, and Ab minor. A summary of the key distributions is provided in Table.2
Table 2: The number of keys in SHTE and Bach datasets
Keys C C#/Db D D#/Eb E F F#/Gb G G#/Ab A A#/Bb B
Major SHTE 6 4 11 8 10 13 1 8 8 11 13 3
Bach 22 - 32 9 10 28 - 59 1 32 25 1
Minor SHTE 9 6 10 - 11 7 17 17 - 9 2 9
Bach 10 - 27 - 20 2 7 44 - 49 2 28
5 Experiments
5.1 Our Model VS Psychological Model
    $PSC_{k}=\sum^{4}_{j=1}\sum^{128}_{r\%12=k}\sum^{N}_{s}c_{k,jrs}$ (14)
For the minor mode(Fig.5(B)(D)), the PSCs and PASWs also play the most important role at tonic(C). However, in the KS model, the scores for mediant(Eb)$KS_{m}inor(Eb)=0.75$ are higher than those for the dominant(G)$KS_{m}inor(G)=0.58$. In contrast, the PSC and PASW values in the SHTE dataset($PSC_{SHTE}(Eb)=0.52$, $PASW_{SHTE}(Eb)=0.45$) and Bach dataset ($PSC_{Bach}(Eb)=0.74$, $PASW_{Bach}(Eb)=0.58$) are lower than those at the dominant(G). This discrepancy can be attributed to the key and necessary role of the dominant chord in the cadences and its frequent use in both datasets. Besides, compared to KS model in SHTE dataset(5(B)), these two metrics for the leading tone(B) show an ascending trend, This is due to the teaching characteristics of the SHTE dataset, which includes more exercises focused on learning chords.
5.2 Generation Evaluation And Analysis
5.2.1 Metrics
To further assess and analyze the quality of the generated music pieces, we employ a comprehensive objective method which is inspired by the method proposed by Yang and Lerch Yang and Lerch (2020). Several features are extracted from the generated samples and two datasets involved in our model. The features include: Pitch Count(PC), the number of different pitches in a music piece; Pitch Class Histogram(PCH), the histogram of 12 pitch classes; Diatonic Pitch Rate(DPR), the proportion of pitches in a music piece that belongs to the key’s diatonic scale; Pitch Range(PR), the pitch range between the highest and the lowest pitch in semitones within a sample; Pitch Interval(PI), average interval between two pitch neighbors, PCTM, the pitch class transition matrix.
5.2.2 Performance Comparison
A) Comparison with datasets
Table. 3 shows the comparison results between the generated samples and the trainning datasets. We randomly select 50 samples of our training datasets SHTE and the J.S. Bach’s(JSB), respectively. Similarly, we randomly select 50 samples of our generated 8-bars long music pieces as the test dataset, covering different keys for testing. We compute the features mentioned above for these three datasets. Both absolute and relative measurements are discussed for these features mentioned above. The absolute measurements specially refer to the mean and the standard deviation, focusing on the characteristics within each dataset. On the other hand, the relative measurements emphasize the features across datasets, with the Kullback–Leibler divergence(KLD) and overlapped area(OA) being employed to evaluate the similarities between the generated music pieces and the two datasets.
SHTE Bach Generated Data
Mean STD Mean STD Mean STD KLD OA KLD OA
PC 8.00 1.79 8.40 2.06 7.70 1.27 0.13 0.67 0.11 0.71
PCH - - - - - - 0.16 0.59 0.21 0.45
DPR 0.96 0.04 0.93 0.04 0.86 0.04 0.02 0.78 0.03 0.83
PR 11.9 3.51 11.7 2.23 9.4 2.2 0.05 0.74 0.01 0.82
PI 2.51 1.01 1.92 0.41 2.33 0.32 0.11 0.55 0.04 0.71
PCTM - - - - - - 0.09 0.45 0.19 0.28
Referring to the evaluation results, the analysis are as follows:
  • Besides, for the PCH and PCTM features,The KLD of PCH is lower for SHTE ($0.16$) than for Bach ($0.21$), meaning the pitch class histogram of the generated data is more aligned with SHTE. The OA values for PCH are $0.59$ (SHTE) and $0.45$ (Bach), also suggest that the generated data resembles SHTE more in pitch class distribution patterns. Similarly, KLD values for PCTM are higher with Bach ($0.19$) than with SHTE ($0.09$), and OA values are $0.45$ (SHTE) and $0.28$ (Bach), suggesting that the pitch transitions of generated data more closely resemble SHTE.
In summary, the generated data aligns more closely with Bach’s melodic characteristics while adopting some unique features from SHTE, such as a tighter pitch class spread and more distinct transition patterns, contributing to a balance of traditional structure and innovative variation.
B) Comparison with Baseline
Given the limited research on music generation using spiking neural networks, we focus on a single existing work based on brain-inspired SNN, referred to as SCSNN Liang and Zeng (2021). Figure 8 provides visual comparison of the mean values of four feature metrics (PC, DPR, PR, and PI) for the SCSNN baseline and our model. The error bars represent the standard deviation for each metric, indicating the variability within the results.
Figure 8: The comparison of the mean values between our model and the SCSNN baseline for the PC, DPR, PR, and PI metrics.
The figure shows that, our model demonstrates a significantly higher mean PC, indicating that it generates music with a broader variety of pitches compared to SCSNN. The slightly higher standard deviation suggests that, while our model produces more varied music, it also exhibits greater variability in pitch diversity. For the DPR feature, our model achieves a higher rate than SCSNN(mean = 0.86 vs. 0.79) and reduced variability (std = 0.04 vs. 0.06), indicating that it generates music with pitches more closely aligned to the key’s diatonic scale, thereby better preserving tonal consistency and producing harmonically coherent samples. Regarding PR, our model significantly outperforms SCSNN in range (mean = 9.4 vs. 5.9) but shows greater variability (std = 2.2 vs. 0.94), suggesting that it generates music with a broader tonal span. This increase in pitch range reflects greater melodic expressiveness and dynamic variety in the generated samples. The higher std could introduce unpredictability, it also highlights the model’s ability to explore a wider range of musical expressions, from simpler, narrower melodies to more expansive, complex ones. Finally, for PI, our model yields a larger average interval (mean = 2.33 vs. 1.68) with slightly lower variability (std = 0.32 vs. 0.38).
Besides, the comparison of two important metrics, KLD and OA, is presented in Fig. 9. Our model achieves lower KLD for PCH, PR, PI, and PCTM, with a particularly notable improvement in DPR, demonstrating better ability of the diatonic pitch structure. Specifically, for DPR, our model outperforms SCSNN with KLD values of 0.02 (SHTE) and 0.03 (Bach) compared to SCSNN’s 0.15 (SHTE) and 0.04 (Bach), indicating better adherence to tonal scales. The higher OA values for DPR further confirm that our model generates music that closely overlaps with the real data distributions, enhancing tonal coherence. For PR, our model shows significantly lower KLD and higher OA values on both datasets compared to SCSNN, indicating that our model’s pitch range better matches the actual data, leading to more realistic melodies. In PCH, our model performs better for SHTE with a KLD of 0.16 compared to SCSNN’s 0.57, while for Bach, both models show similar KLD values (0.21 for our model and 0.20 for SCSNN). However, SCSNN achieves higher OA for Bach, suggesting slightly better alignment for that dataset. Regarding PC, although both models have similar KLD values, our model achieves higher OA (0.67 for SHTE and 0.71 for Bach), indicating more consistent pitch generation. For PI, while SCSNN performs better for SHTE with a lower KLD (0.07 vs. 0.11), our model excels for Bach with a significantly lower KLD of 0.04 versus SCSNN’s 0.41. Combined with higher OA values, this result highlights our model’s improved accuracy in capturing pitch transitions. Lastly, in PCTM, our model outperforms SCSNN across both datasets, with lower KLD values (0.09 for SHTE and 0.19 for Bach) and higher OA, indicating better modeling of pitch class transitions.
Figure 9: The performance comparison of KLDs and OAs for the features between our model and the baseline.
Overall, the results further indicate that the proposed model effectively captures both the tonality characteristics and the melody adaptability needed for generating diverse and musical content. The model’s ability to blend well-established tonal principles with novel musical expressions underscores its potential for generating musically meaningful compositions across various styles and datasets. This capability is essential for our future work on harmonic learning, emotion perception, and stylistic composition.
@@ 7 Acknowledgments
7.1 Author contributions
Q.L designed the study, implemented the code, conducted the experiments, and drafted the manuscript. Y.Z. supervised the project and contributed to the manuscript writing. MHR.T. assisted with code maintenance.
7.2 Declaration of interests